# neutron star; dark matter; gravitational waves; tidal deformability; neutron decay

### v2

@@ 1 Introduction
Compact objects represent a unique class of astrophysical bodies characterized by incredibly high densities and strong gravitational fields. This group includes white dwarfs (WDs), neutron stars (NSs), and black holes (BHs), each formed through the evolutionary endpoints of stellar collapse and distinguished by the degeneracy pressure mechanisms or gravitational forces that prevent further collapse. WDs are sustained by electron degeneracy pressure, limiting their masses to the Chandrasekhar limit of approximately $1.4$ $M_{\mathrm{\odot}}\,$handrasekhar (1939), while NSs are held up by neutron degeneracy pressure and strong nuclear forces Lattimer and Prakash (2004); Baym et al. (2018). BHs, by contrast, lack any such internal support mechanism and represent the ultimate state of stellar evolution, where collapse leads to singularities shielded by event horizons Shapiro and Teukolsky (1983).
Among compact objects, NSs represent ideal environments to probe extreme states of matter, gravitational physics, and, potentially, DM interactions Silk et al. (2010). Primarily composed of densely packed neutrons, NSs have masses ranging from about 1.4 to 2.5 $M_{\mathrm{\odot}}\,$attimer and Prakash (2004); Antoniadis et al. (2013); Özel and Freire (2016), compressed into a radius of roughly 10–12 km Lattimer and Prakash (2007); Steiner et al. (2010); Oertel et al. (2017). The central densities of NSs can exceed several times the nuclear saturation density ($\rho_{\mathrm{nucl}}\sim 2.7\times 10^{14}\,\mathrm{g\,cm^{-3}}$), and can reach values as high as $5\times 10^{15}\,\mathrm{g\,cm^{-3}}$ (Figure 4 in Lattimer and Prakash (2011)) close to the core, orders of magnitude greater than the densities found in ordinary atomic nuclei Lattimer and Prakash (2016); Baym et al. (2018). Such extreme densities lead to intense gravitational fields, with surface gravity
@@
[…]
@@
[…]
@@
DM could also leave signatures in the GW signals emitted by BNS mergers. During the inspiral and merger phases of such collisions, the presence of DM could alter the tidal deformability ($\Lambda$) and the gravitational waveform, potentially leading to detectable deviations from the predictions of purely baryonic models Ellis et al. (2018); Kopp et al. (2018); Bhattacharya et al. (2023). This idea gained traction following the detection of GWs from GW170817. Remarkably, the GW190814 Abbott et al. (2020) event involved a $\sim$23232323 $M_{\mathrm{\odot}}\,$ and a compact object with a mass that was challenging to classify: either a high-mass NS or a low-mass BH Most et al. (2020). The uncertainty about the nature of the lighter companion led to speculation that it might be an NS with a significant fraction of DM Das et al. (2021). Future observations by both current detectors and incoming interferometers (such as the Einstein Telescope Hild et al. (2011); Branchesi et al. (2023) and the Cosmic Explorer Abbott et al. (2017); Evans et al. (2023)) could provide crucial insights into the existence of DM within NSs. Finally, ongoing X-ray observations by the NICER mission Gendreau et al. (2012), which aims to measure NS masses and radii with great precision, offer another method for detecting DM effects by looking for any deviations from the mass–radius relationship predicted by purely baryonic frameworks.
@@ 2 Dark Matter Capture and Accumulation
One of the primary mechanisms for DM accumulation within NSs is gravitational capture. As stars orbit around the center of a galaxy, they encounter significant fluxes of DM particles. Interactions between DM and SM particles within stars can lead to energy loss, potentially trapping DM in the star’s gravitational field. This capture process would be particularly efficient due to the extreme density of NSs. Assuming a typical NS of $M\sim 1.5$ $M_{\mathrm{\odot}}\,$and $R\sim 12$ km, the corresponding escape speed at its surface is $v_{\mathrm{esc}}=\sqrt{2M/R}\approx 0.6$, as measured by an observer at rest on the surface of the NS.
@@
$\sigma_{\mathrm{cap}}=\begin{cases}\sigma_{0}\left(\frac{m_{\mathrm{n}}}{m_{\chi}}\right),&\mathrm{if\;}m_{\mathrm{evap}}<m_{\chi}<m_{\mathrm{n}}\\ \sigma_{0},&\mathrm{if\;}m_{\mathrm{n}}\leq m_{\chi}\leq\mathrm{PeV}\\ \sigma_{0}\left(\frac{m_{\chi}}{\mathrm{PeV}}\right),&\mathrm{if\;}m_{\chi}>\mathrm{PeV}\,,\end{cases}$ (3)
@@
where $\sigma=\frac{\pi^{2}k_{\mathrm{B}}^{4}}{60\hbar^{3}c^{2}}=5.670400\times 10^{-8}\;\mathrm{W\,m^{-2}\,K^{-4}}$ is the Stefan–Boltzmann constant (with $k_{\mathrm{B}}=1.380\;\mathrm{J\,K^{-1}}$ being the Boltzmann constant) and $g=GM/R$ accounts for gravitational effects and plays the role of an efficiency factor representing the fraction of the rest-mass energy converted into radiation during the accretion process  Ikhsanov (2003). The second equality, in which $\dot{M}$ is the rate at which mass is added to the star (Equation (2)), holds by assuming that the NS behaves as a black-body radiator, whose luminosity $L=A\sigma T_{\mathrm{s}}^{4}$ follows the Stefan–Boltzmann law Yin et al. (2011); Carroll and Ostlie (2017) for an approximately spherical NS of surface area $A=4\pi R^{2}$. The maximum photon luminosity $L_{\mathrm{max}}$ is obtained when $P_{\sigma_{\chi}}\approx 1$ and $P_{\sigma_{\mathrm{v}}}\approx 1$. Under these conditions, $L_{\mathrm{max}}$ can be computed close to the Solar System ($\sim$10101010 pc) as $L_{\mathrm{max}}\approx 7\times 10^{24}\,\mathrm{GeV/s}$. Remarkably, for an ordinary NS with $R\sim 10$ km, this value corresponds to a superficial, effective temperature of $\sim$2000200020002000 K that may be detectable by cutting-edge infrared telescopes Chatterjee et al. (2023) such as the James Webb Space Telescope Gardner et al. (2006); Kalirai (2018); Rigby et al. (2023).
To investigate the cooling profile, previous works Yakovlev et al. (1999, 2001); Yakovlev and Pethick (2004) have relied on the “minimal cooling” assumption, meaning that nucleons and charged leptons remain in $\beta$-equilibrium throughout the NS’s evolution, with no internal mechanisms contributing to its heating ($L_{\mathrm{int}}^{\infty}=0$). The resulting $T_{s}$ of isolated, old NSs was estimated to be $<$$1000$ K. However, when DM annihilation is included ($L_{\mathrm{DM}}^{\infty}\neq 0$), the predicted surface temperatures rise to $T_{s}\simeq(2-3)\times 10^{3}$ K. This suggests that the measurements of $T_{s}$ of old NSs could provide critical insights and constraints for DM models.
@@
[…]
@@ 2.1 Self-Annihilating Dark Matter
×$\times\left(1-\frac{1-e^{-A^{2}}}{A^{2}}\right)\left(\frac{v_{\mathrm{esc}}}{1.9\times 10^{5}\;\mathrm{km\,s^{-1}}}\right)^{2}\left(\frac{220\,\mathrm{km\,s^{-1}}}{v_{\chi}}\right)^{2}\,,$ (9)
@@
where $\tau$ is the average lifetime of an NS (with mass $\sim$1.51.51.51.5 $M_{\mathrm{\odot}}\,$and radius $\sim$12121212 km) and $C_{\mathrm{max}}$ the maximum capture rate.
@@ 2.2 Black Hole Formation
[…]
@@
    $M_{i}\equiv\begin{cases}M_{f}\sim M_{\mathrm{pl}}^{3}/m_{\chi}^{2}\sim\,M_{\odot}\left(\frac{\mathrm{GeV^{2}}}{m_{\chi}^{2}}\right)\,\;\;&\mathrm{for\;\;fermionic\;\;DM}\,,\\ M_{b}\sim\frac{2M_{\mathrm{pl}}^{2}}{m_{\chi}}\left(1+\frac{\lambda}{32\pi}\frac{M_{\mathrm{pl}}^{2}}{m_{\chi}^{2}}\right)^{1/2}\;\;&\mathrm{for\;\;bosonic\;\;DM}\,,\end{cases}$ (16)
@@
[…]
@@ 3.4 Mass–Radius and Tidal–Mass Relations
While the masses of the DM candidates have been considered fixed so far, the effects of a variable $m_{\chi}$ have been demonstrated by Routaray et al. Routaray et al. (2023). Their findings indicate that lighter DM particles ($\sim$300300300300 MeV) are more likely to form extended DM halos, whereas heavier DM candidates tend to be confined to the core of DANSs. Furthermore, they stressed that with intermediate DM mass ($\sim$500500500500 MeV) cores form only with high $f_{\mathrm{DM}}$.
[…]
Such a trend is confirmed even when considering different EoSs for BM. For instance, Ref. Liu et al. (2023) derives the BM EoS from a Brueckner–Hartree–Fock approach, covering all $f_{\mathrm{DM}}\in[0,1]$. Their findings confirm that massive DM particles, low fractions, and weak couplings favor core formation and, in turn, result in smaller and denser DANSs. Additionally, they found a couple of interesting outcomes. First, stable high-$f_{\mathrm{DM}}$ stars can be produced under specific conditions: the hosting NSs must be very light (with mass $\leq 0.5$ $M_{\mathrm{\odot}}\,$if $f_{\mathrm{DM}}\sim 0.9$—see Figure 3 in Liu et al. (2023)) and DM should accumulate through exotic mechanisms such as accretion of normal matter onto a dark star Ciarcelluti and Sandin (2011); Ellis et al. (2018); Gleason et al. (2022). Second, twin DANSs may exist, namely, a low-$f_{\mathrm{DM}}$ star with a DM core and a high-$f_{\mathrm{DM}}$ star with a DM halo both with $M_{\mathrm{DANS}}\sim 0.5$ $M_{\mathrm{\odot}}\,$and $R_{\mathrm{DANS}}\sim 12$ km. These twins would be indistinguishable based solely on mass and optical radius measurements, yet their internal structure and related properties would be fundamentally different (due to the way DM is distributed).
@@
The (dimensionless) tidal deformability $\Lambda$ is computed by means of Equation (23), where $M\equiv M_{\mathrm{DANS}}$ and $R\equiv R_{\mathrm{DANS}}$, i.e., the outermost radius. As for the mass–radius relations, in Figure 5 we report how a nuclear EoS (as BSk22) is modified if DM accumulates within the star according to the afore-mentioned model of fermionic, interacting DM. For a given total gravitational mass, the presence of DM condensed in a core results in a smaller tidal deformability Ivanytskyi et al. (2020); Liu et al. (2024) compared to a pure NS. When a DM halo forms, both the mass and radius of the NS decrease. However, in the scaling relation $\Lambda\propto(M_{\mathrm{DANS}}/R_{\mathrm{DANS}})^{-5}$ (Equation (23)), the reduction in radius has a dominant effect over the reduction in mass, resulting in an overall decrease in tidal deformability. From the perspective of a distant observer, these changes would be interpreted as an effective softening of the EoS. Indeed, with a softer EoS less pressure can be exerted to balance the gravitational self-interaction, so the star becomes more compact with a smaller radius. This increased compactness results in the star being more resistant to deformation, thus lowering its $\Lambda$. However, as pointed out in Giangrandi et al. (2023), this effect may be not easy to identify. As for the M–R curves, the modifications in the $\Lambda$–M relations alone are not sufficient to establish the presence of DM within NSs. Indeed, both the DM capture and the presence of an additional degree of freedom (such as hyperons) can contribute to the softening of the EoS. As a consequence, distinguishing these scenarios based on the measurement of $\Lambda$ and/or the observed radius mass would be challenging. In other words, despite completely different internal structures, the same mass and radius could correspond to two distinct situations: a DANS with a certain fraction of DM and a pure NS containing hyperons, SQM, or other new particles. In contrast, the existence of a DM halo leads to a substantial rise in the tidal deformability because of the significant increase in $R_{\mathrm{DANS}}$.
As for the dependence on the coupling, $\Lambda$ increases monotonically with larger $g_{v}$, eventually favoring the formation of a DM halo. Das et al. Das et al. (2022) have emphasized this evidence is important in differentiating a single-fluid model from a two-fluid approach when describing DANSs. Indeed, from the analysis in Das et al. (2019) (where DM is coupled to BM through a Higgs portal interaction, forming one fluid), the stellar properties, including $\Lambda$, always decrease in the presence of DM. In contrast, in the two-fluid scenario, $\Lambda$ can rise beyond the observational limits when a halo extends over $\sim$20202020 km. Due to such a great radius, the resulting curves do not usually match experimental observations, making configurations with a halo disfavored by these DM models.

### v3

@@ 1 Introduction
Compact objects represent a unique class of astrophysical bodies characterized by incredibly high densities and strong gravitational fields. This group includes white dwarfs (WDs), neutron stars (NSs), and black holes (BHs), each formed through the evolutionary endpoints of stellar collapse and distinguished by the degeneracy pressure mechanisms or gravitational forces that prevent further collapse. WDs are sustained by electron degeneracy pressure, limiting their masses to the Chandrasekhar limit of approximately $1.4$ $M_{\mathrm{\odot}}\,$Chandrasekhar (1939), while NSs are held up by neutron degeneracy pressure and strong nuclear forces Lattimer and Prakash (2004); Baym et al. (2018). BHs, by contrast, lack any such internal support mechanism and represent the ultimate state of stellar evolution, where collapse leads to singularities shielded by event horizons Shapiro and Teukolsky (1983).
Among compact objects, NSs represent ideal environments to probe extreme states of matter, gravitational physics, and, potentially, DM interactions Silk et al. (2010). Primarily composed of densely packed neutrons, NSs have masses ranging from about 1.4 to 2.5 $M_{\mathrm{\odot}}\,$Lattimer and Prakash (2004); Antoniadis et al. (2013); Özel and Freire (2016), compressed into a radius of roughly 10–12 km Lattimer and Prakash (2007); Steiner et al. (2010); Oertel et al. (2017). The central densities of NSs can exceed several times the nuclear saturation density ($\rho_{\mathrm{nucl}}\sim 2.7\times 10^{14}\,\mathrm{g\,cm^{-3}}$), and can reach values as high as $5\times 10^{15}\,\mathrm{g\,cm^{-3}}$ (Figure 4 in Lattimer and Prakash (2011)) close to the core, orders of magnitude greater than the densities found in ordinary atomic nuclei Lattimer and Prakash (2016); Baym et al. (2018). Such extreme densities lead to intense gravitational fields, with surface gravity
@@
The intense gravitational collapse leading to the formation of an NS heats its interior to a temperature of $\sim$$10\,\mathrm{MeV}\simeq 10^{11}$ K Yakovlev and Urpin (1980); Yakovlev and Pethick (2004); Weber et al. (2007). However, NSs cool rapidly over time via neutrino emission, stabilizing to surface temperatures in the range of $10^{6}$$10^{8}$ K after a few million years Pons et al. (1999); Ho et al. (2007); Potekhin et al. (2015). Consequently, $T$ drops significantly, becoming about 3–4 orders of magnitude lower. That is why NSs are considered cold compact objects, for which the temperature, when expressed in MeV, is approximately zero. Despite this cooling, thermal and dynamic processes remain active and can provide clues about the underlying microphysics of dense matter.
@@
Early explanations focused on Massive Astrophysical Compact Halo Objects (MACHOs)—such as brown dwarfs, BHs, and faint stars Lasserre (2000)—or on modifications to gravitational laws, like the modified Newtonian dynamics (MOND) theory Famaey and McGaugh (2012). However, the inability of these hypotheses to match the results from microlensing experiments Tisserand et al. (2007) and observations within the Bullet Cluster Clowe et al. (2006) led to the conclusion that a new type of matter (i.e., particles) was needed to explain cosmic structures. Standard Model (SM) particles like neutrinos were also discarded Boyarsky et al. (2019), so the leading candidate became either weakly interacting massive particles (WIMPs) Pagels and Primack (1982); Jungman et al. (1996); Bertone et al. (2005), interacting via the weak nuclear force and gravity, or axions, lightweight particles originally proposed to solve the strong CP problem in quantum chromodynamics Weinberg (1972); Peccei and Quinn (1977). Overall, several theories have been proposed throughout the decades to explain the nature of DM, spanning a mass range of more than 90 orders of magnitude Cirelli et al. (2024); namely, from $\sim$$10^{-22}$ eV (fuzzy DM Hu et al. (2000); Hui et al. (2017); Ferreira (2021)) to $10^{2}$ $M_{\mathrm{\odot}}\,$(primordial BHs Carr et al. (2021); Carr and Kuhnel (2020)).
@@
DM could also leave signatures in the GW signals emitted by BNS mergers. During the inspiral and merger phases of such collisions, the presence of DM could alter the tidal deformability ($\Lambda$) and the gravitational waveform, potentially leading to detectable deviations from the predictions of purely baryonic models Ellis et al. (2018); Kopp et al. (2018); Bhattacharya et al. (2023). This idea gained traction following the detection of GWs from GW170817. Remarkably, the GW190814 Abbott et al. (2020) event involved a $\sim$$23$ $M_{\mathrm{\odot}}\,$BH and a compact object with a mass that was challenging to classify: either a high-mass NS or a low-mass BH Most et al. (2020). The uncertainty about the nature of the lighter companion led to speculation that it might be an NS with a significant fraction of DM Das et al. (2021). Future observations by both current detectors and incoming interferometers (such as the Einstein Telescope Hild et al. (2011); Branchesi et al. (2023) and the Cosmic Explorer Abbott et al. (2017); Evans et al. (2023)) could provide crucial insights into the existence of DM within NSs. Finally, ongoing X-ray observations by the NICER mission Gendreau et al. (2012), which aims to measure NS masses and radii with great precision, offer another method for detecting DM effects by looking for any deviations from the mass–radius relationship predicted by purely baryonic frameworks.
@@ 2 Dark Matter Capture and Accumulation
One of the primary mechanisms for DM accumulation within NSs is gravitational capture. As stars orbit around the center of a galaxy, they encounter significant fluxes of DM particles. Interactions between DM and SM particles within stars can lead to energy loss, potentially trapping DM in the star’s gravitational field. This capture process would be particularly efficient due to the extreme density of NSs. Assuming a typical NS of $\sim 1.5 MM_{\mathrm{\odot}}\,$and $R\sim 12$ km, the corresponding escape speed at its surface is $v_{\mathrm{esc}}=\sqrt{2M/R}\approx 0.6$, as measured by an observer at rest on the surface of the NS.
@@
[…]
@@
where $\sigma=\frac{\pi^{2}k_{\mathrm{B}}^{4}}{60\hbar^{3}c^{2}}=5.670400\times 10^{-8}\;\mathrm{W\,m^{-2}\,K^{-4}}$ is the Stefan–Boltzmann constant (with $k_{\mathrm{B}}=1.380\;\mathrm{J\,K^{-1}}$ being the Boltzmann constant) and $g=GM/R$ accounts for gravitational effects and plays the role of an efficiency factor representing the fraction of the rest-mass energy converted into radiation during the accretion process  Ikhsanov (2003). The second equality, in which $\dot{M}$ is the rate at which mass is added to the star (Equation (2)), holds by assuming that the NS behaves as a black-body radiator, whose luminosity $L=A\sigma T_{\mathrm{s}}^{4}$ follows the Stefan–Boltzmann law Yin et al. (2011); Carroll and Ostlie (2017) for an approximately spherical NS of surface area $A=4\pi R^{2}$. The maximum photon luminosity $L_{\mathrm{max}}$ is obtained when $P_{\sigma_{\chi}}\approx 1$ and $P_{\sigma_{\mathrm{v}}}\approx 1$. Under these conditions, $L_{\mathrm{max}}$ can be computed close to the Solar System ($\sim$$10$ pc) as $L_{\mathrm{max}}\approx 7\times 10^{24}\,\mathrm{GeV/s}$. Remarkably, for an ordinary NS with $R\sim 10$ km, this value corresponds to a superficial, effective temperature of $\sim$$2000$ K that may be detectable by cutting-edge infrared telescopes Chatterjee et al. (2023) such as the James Webb Space Telescope Gardner et al. (2006); Kalirai (2018); Rigby et al. (2023).
To investigate the cooling profile, previous works Yakovlev et al. (1999, 2001); Yakovlev and Pethick (2004) have relied on the “minimal cooling” assumption, meaning that nucleons and charged leptons remain in $\beta$-equilibrium throughout the NS’s evolution, with no internal mechanisms contributing to its heating ($L_{\mathrm{int}}^{\infty}=0$). The resulting $T_{s}$ of isolated, old NSs was estimated to be <<$1000$ K. However, when DM annihilation is included ($L_{\mathrm{DM}}^{\infty}\neq 0$), the predicted surface temperatures rise to $T_{s}\simeq(2-3)\times 10^{3}$ K. This suggests that the measurements of $T_{s}$ of old NSs could provide critical insights and constraints for DM models.
@@
Such an imbalance provides a supplementary heating mechanism, known as rotochemical heating. When included Fernandez and Reisenegger (2005); Villain and Haensel (2005); Petrovich and Reisenegger (2010); González-Jiménez et al. (2015); Yanagi et al. (2020) ($L_{\mathrm{int}}^{\infty}=L_{\mathrm{rc}}^{\infty}\neq 0$), it can raise $T_{s}$ up to $10^{6}$ K, as confirmed by observations Kargaltsev et al. (2004); Rangelov et al. (2017); Pavlov et al. (2017); Gonzalez-Caniulef et al. (2019). This makes the interpretation of NS cooling profiles and surface temperature measurements more challenging, as the intrinsic heating of the NS is now degenerate with the DM contribution. Interestingly, Hamaguchi et al. Hamaguchi et al. (2019) have demonstrated that this degeneracy can be partially broken: the DM heating effect can still be observed if the initial period of ordinary NSs is relatively large, though we need to improve the current knowledge on nucleon pairing gaps as well as to evaluate the initial period of the pulsars more accurately. Furthermore, the lower limit $T_{s}\gtrsim(2$$3)\times 10^{3}$ K should still hold because the rotochemical mechanism may only increase the surface temperature. Hence, observing an NS with much lower $T_{s}$ would ”exclude the DM heating, and severely constrain DM models”.
@@ 2.1 Self-Annihilating Dark Matter
$\times\left(1-\frac{1-e^{-A^{2}}}{A^{2}}\right)\left(\frac{v_{\mathrm{esc}}}{1.9\times 10^{5}\;\mathrm{km\,s^{-1}}}\right)^{2}\left(\frac{220\,\mathrm{km\,s^{-1}}}{v_{\chi}}\right)^{2}\,$, (9)
@@
where $\tau$ is the average lifetime of an NS (with mass $\sim$$1.5$ $M_{\mathrm{\odot}}\,$and radius $\sim$$12$ km) and $C_{\mathrm{max}}$ the maximum capture rate.
@@ 2.2 Black Hole Formation
    computed for an average NS with core density $\rho_{\mathrm{NS}}\sim 10^{15}\,\mathrm{g/cm^{3}}$ and absolute temperature $\sim$$10^{8}$ K.
@@
[…]
@@
Figure 2: Left panel: Bounds on the (heavy) fermionic DM–nucleon cross-section obtained from the observation of old pulsars in the Milky Way that have not been converted to BHs. Ref. Bramante et al. (2018) (from which the plot is taken) also reported constraints from terrestrial direct experiments such as Xenon1T Aprile et al. (2019a, b, c, 2021) and xenon neutrino floor Ruppin et al. (2014), as well as potential observations of binary NS mergers accompanied by kilonovae (with localization precision up to 1 kpc within Milky Way-like spiral galaxies). Right panel: Bounds on DM-neutron cross-section obtained with the gravitational collapse condition applied to an old NS with a thermalized core of bosonic DM when a BEC is not formed. The representative NS is assumed to have an average lifetime of $10^{10}$ years and a mean temperature of $10^{5}$ K, according to Garani et al.’s Garani et al. (2019) model, from which the plot is adapted. In the red areas (corresponding to different values of $\rho_{\chi}$), the accumulated DM triggers the collapse to a BH. In the green areas, DM fails to thermalize with neutrons. The hatched regions indicate where the BH evaporates before it can destroy the NS, thereby weakening the constraints. The Xenon1T limit on DM–neutron cross-sections is represented by the dark blue-shaded region. The updated upper limits from the Lux-Zeplin 2022 Aalbers et al. (2023) and 2024 Aalbers et al. (2024) campaigns are reported in light blue. Note that these experiments explored the WIMP range up to $\sim$$10^{4}$ GeV (see also Figure 3 below). Remarkably, in Garani et al. (2019) similar analyses have been performed for DM scattering with protons and muons as well as when considering bosonic DM that forms a BEC or fermionic DM.
@@ 3.3 Fermionic DM
As follows from the EoSs in Eqs. (30)-(31) and Eqs. (36)-(37), the phenomenology of DANSs is largely controlled by a set of poorly known microscopic parameters, whose estimation requires a dedicated statistical framework. In recent years, Bayesian inference has become a widely adopted approach to constrain NS models using astrophysical observations, and has been extensively applied to infer the properties of dense nuclear matter from mass, radius, and tidal-deformability measurements Huang et al. (2024); Grundler et al. (2025); Imam et al. (2025); Cartaxo et al. (2025).
The same framework is even more relevant for bosonic/fermionic DM, whose parameters – such as the DM candidate mass and the strength of its interactions – are far less constrained than their baryonic counterparts, and can span wide ranges without being excluded by current observations.
In this context, theoretical information on the microscopic parameters is combined with astrophysical data within the Bayesian inference through Bayes’ theorem
$p\left(\vec{\theta};\vec{x}\right)=\frac{\pi\left(\vec{\theta}\right)\times\mathcal{L}\left(\vec{x};\vec{\theta}\right)}{p\left(\vec{x}\right)}\;,$ (38)
where $\vec{\theta}$ and $\vec{x}$ denote the set of model parameters and observational data, respectively. In Eq. (38), $\pi(\vec{\theta})$ and $\mathcal{L}(\vec{x};\vec{\theta})$ represent the prior distribution and the likelihood function, while $p(\vec{x})$, known as evidence, serves only as a normalization.
Therefore, the combination of available data, such as GW signals from the LIGO-Virgo collaboration and NICER constraints on the mass and the radius of NSs, allows to construct posterior distributions for the DM couplings and masses. This strategy has already proven effective in recent studies of bosonic and fermionic DANSs Grippa et al. (2024); Liu et al. (2025); Arvikar et al. (2025); Santos et al. (2025), and is expected to play an increasingly important role as observational constraints improve.
@@ 3.4 Mass–Radius and Tidal–Mass Relations
While the masses of the DM candidates have been considered fixed so far, the effects of a variable $m_{\chi}$ have been demonstrated by Routaray et al. Routaray et al. (2023). Their findings indicate that lighter DM particles ($\sim$$300$ MeV) are more likely to form extended DM halos, whereas heavier DM candidates tend to be confined to the core of DANSs. Furthermore, they stressed that with intermediate DM mass ($\sim$$500$ MeV) cores form only with high $f_{\mathrm{DM}}$.
Regarding bosonic DM, as highlighted by Ellis et al. Ellis et al. (2018), once a positive pressure is established for the DM component to prevent the formation of a BH at the center of the NS, both bosonic and fermionic DM produce similar effects on DANSs. This was found in Giangrandi et al. (2023), where bosonic DM is coupled with nuclear EoSs as IST, DD2, and $\mathrm{DD2}\Lambda$ (the latter accounting for the presence of hyperons that soften the EoS). In all three cases, when considering a DM candidate with mass $1$ GeV and relatively small DM fractions $f_{\mathrm{DM}}$ (1%, 3%, and 5%), a DM core is formed within the NS that leads to a decrease in both the mass and the radius. As $f_{\mathrm{DM}}$ increases, the core becomes heavier, and $M_{\mathrm{DANS}}$ and $R_{\mathrm{DANS}}$ decrease further. Thus, the presence of a DM core effectively softens the baryonic EoS. However, if too much DM is accumulated near the core and/or light DM candidates are considered ($\sim$$100$ MeV), a core no longer exists; rather, a halo forms due to the weaker gravitational interaction.
Such a trend is confirmed even when considering different EoSs for BM. For instance, Ref. Liu et al. (2023) derives the BM EoS from a Brueckner–Hartree–Fock approach, covering all $f_{\mathrm{DM}}\in[0,1]$. Their findings confirm that massive DM particles, low fractions, and weak couplings favor core formation and, in turn, result in smaller and denser DANSs. Additionally, they found a couple of interesting outcomes. First, stable high-$f_{\mathrm{DM}}$ stars can be produced under specific conditions: the hosting NSs must be very light (with mass $\leq 0.5$ $M_{\mathrm{\odot}}\,$if $f_{\mathrm{DM}}\sim 0.9$—see Figure 3 in Liu et al. (2023)) and DM should accumulate through exotic mechanisms such as accretion of normal matter onto a dark star Ciarcelluti and Sandin (2011); Ellis et al. (2018); Gleason et al. (2022). Second, twin DANSs may exist, namely, a low-$f_{\mathrm{DM}}$ star with a DM core and a high-$f_{\mathrm{DM}}$ star with a DM halo both with $_{\mathrm{DANS}}\sim 0.5 MM_{\mathrm{\odot}}\,$and $R_{\mathrm{DANS}}\sim 12$ km. These twins would be indistinguishable based solely on mass and optical radius measurements, yet their internal structure and related properties would be fundamentally different (due to the way DM is distributed).
@@
The (dimensionless) tidal deformability $\Lambda$ is computed by means of Equation (23), where $M\equiv M_{\mathrm{DANS}}$ and $R\equiv R_{\mathrm{DANS}}$, i.e., the outermost radius. As for the mass–radius relations, in Figure 5 we report how a nuclear EoS (as BSk22) is modified if DM accumulates within the star according to the afore-mentioned model of fermionic, interacting DM. For a given total gravitational mass, the presence of DM condensed in a core results in a smaller tidal deformability Ivanytskyi et al. (2020); Liu et al. (2024) compared to a pure NS. When a DM core forms, both the mass and radius of the NS decrease. However, in the scaling relation $\Lambda\propto(M_{\mathrm{DANS}}/R_{\mathrm{DANS}})^{-5}$ (Equation (23)), the reduction in radius has a dominant effect over the reduction in mass, resulting in an overall decrease in tidal deformability. From the perspective of a distant observer, these changes would be interpreted as an effective softening of the EoS. Indeed, with a softer EoS less pressure can be exerted to balance the gravitational self-interaction, so the star becomes more compact with a smaller radius. This increased compactness results in the star being more resistant to deformation, thus lowering its $\Lambda$. However, as pointed out in Giangrandi et al. (2023), this effect may be not easy to identify. As for the M–R curves, the modifications in the $\Lambda$–M relations alone are not sufficient to establish the presence of DM within NSs. Indeed, both the DM capture and the presence of an additional degree of freedom (such as hyperons) can contribute to the softening of the EoS. As a consequence, distinguishing these scenarios based on the measurement of $\Lambda$ and/or the observed radius mass would be challenging. In other words, despite completely different internal structures, the same mass and radius could correspond to two distinct situations: a DANS with a certain fraction of DM and a pure NS containing hyperons, SQM, or other new particles. In contrast, the existence of a DM halo leads to a substantial rise in the tidal deformability because of the significant increase in $R_{\mathrm{DANS}}$.
As for the dependence on the coupling, $\Lambda$ increases monotonically with larger $g_{v}$, eventually favoring the formation of a DM halo. Das et al. Das et al. (2022) have emphasized this evidence is important in differentiating a single-fluid model from a two-fluid approach when describing DANSs. Indeed, from the analysis in Das et al. (2019) (where DM is coupled to BM through a Higgs portal interaction, forming one fluid), the stellar properties, including $\Lambda$, always decrease in the presence of DM. In contrast, in the two-fluid scenario, $\Lambda$ can rise beyond the observational limits when a halo extends over $\sim$2020 km. Due to such a great radius, the resulting curves do not usually match experimental observations, making configurations with a halo disfavored by these DM models.